Adding vectors Algebraically: Find the vector components along the directions noted in parenthesis.

1. a car displaced 45°north of east by 10.0km (north and east)
1
Expert's answer
2012-10-15T12:06:28-0400
Adding vectors Algebraically: Find the vector components along the directions noted in parenthesis.
&
1. a car displaced 45њnorth of east by 10.0km (north and east)

The car's north and east components are

N = 10[km]*sin(π/4) = 10/√2 [km],

E = 10[km]*cos(π/4) = 10/√2 [km].

So, the car's coordinates are (10/√2, 10/√2).